# Probabilistic Soundness Guarantees in LLM Reasoning Chains

### v1

@@ 4.4 RQ4: Is ARES useful for selecting Best-of-N generations?
Results in Table 3 show that when using the score of the final step—a stricter and often more decisive measure—ARES significantly outperforms all other methods. Notably, the performance of simpler methods like Entail-Prev collapses on this stricter metric. This highlights ARES’s unique strength in maintaining a sound evaluation of steps throughout the entire reasoning chain, making its final assessment particularly reliable. Therefore, ARES is a strong and robust predictor of downstream task performance.
4.5 Ablations
@@
Inserting Irrelevant Claims.
Table A9 shows that for reasoning chains about trees of depth 5 with 3 sources, other methods show more significant performance drops while ARES maintains high performance. For wide reasoning trees, other methods don’t drop as much, but in trees with greater depth, their error rates increase significantly while ARES remains stable.
Inserting Benign Errors.
We also examined cases where non-existing rules are inserted in reasoning chains but don’t affect later steps. Table A10 shows that all methods perform equally well when errors don’t cause downstream propagation.
We allow using different for the flexibility to not include all the base claims in the premise, and we want to see the impact of different design choices. Table A8 shows ablations for using vs. as well as using granular vs binary entailment models (which use strict ). consistently performs better with a probabilistic entailment model, while a binary entailment model sometimes benefits from . Thus we can effectively choose , including all base claims, which can greatly reduce the actual computation cost as the effective sample size now only depends on entailment of derived claims.
@@ 4.6 Discussion of Errors
Our inspection of the data and error detection outputs reveals some insights. Entail-Base fails on PRMBench because judging entailment in long math derivations is challenging. Both LLM-Judge and Entail-Base fail in DeltaBench, with Entail-Base struggling to judge entailment in very long reasoning chains. In naturally occurring datasets, error propagation is limited and not always annotated, so Entail-Prev performs close to ARES. However, synthetic data shows Entail-Prev fails with propagated errors. LLM-Judge sometimes fails to follow instructions, outputting incorrect numbers of scores relative to claims being judged. Pairwise methods in ROSCOE and ReCEval depen cannot detect complex errors that need multiple claims as premise. ARES can only improve upon entailment models that can already do correct entailment.
@@ 5 Related Work
Reasoning Chain Verifiers.
Primary approaches to verifying reasoning chains include using LLM Judges (Tyagi et al., 2024; He et al., 2024, 2025) or Process Reward Models (PRMs) (Lightman et al., 2023). More recent verifiers incorporate logic but have limitations: ROSCOE (Golovneva et al., 2023) and ReCEval (Prasad et al., 2023) often rely on pairwise contradiction, which struggles with complex premises, while PARC (Mukherjee et al., 2025) offers only a binary classification of soundness. In contrast, our work introduces a probabilistic framework that allows for a more nuanced assessment of each claim.
Evaluating Reasoning Error Detectors.
Numerous benchmarks evaluate CoT reasoning error detectors via step-by-step annotations, such as GridPuzzle (Tyagi et al., 2024), REVEAL (Jacovi et al., 2024), PRMBench (Song et al., 2025), ProcessBench (Zheng et al., 2024), and DeltaBench (He et al., 2025). However, they lack a consistent definition of error. To establish a clear standard, our work adopts a unified definition of soundness, encompassing concepts like validity and groundedness from Lee and Hockenmaier (2025) as well as propagated error from Mukherjee et al. (2025). We define a step as unsound if it is not logically entailed by the set of correct preceding claims. Because existing benchmarks do not uniformly apply this standard, we create synthetic datasets for robust evaluation.
Probabilistic Guarantees.
Our work joins a trend of using statistical guarantees for reliability, common in high-stakes domains (Fayyad et al., 2024; McShane et al., 2023; Lindemann et al., 2023) and in explainable AI for certifying the stability of feature attributions (Jin et al., 2025). We apply this principle to natural language reasoning by providing soundness guarantees over a reasoning chain’s entire directed acyclic graph (DAG) structure. This contrasts with frameworks like BIRD (Feng et al., 2025), which calibrate individual components such as the entailment model, whereas we provide formal guarantees for the multi-step process itself. Our model-agnostic framework is therefore complementary and can leverage a better-calibrated model to make its certifications more reliable.
@@ 6 Conclusion
Current methods cannot reliably detect LLM reasoning errors that propagate. To overcome these limitations, we introduce Autoregressive Reasoning Entailment Stability (ARES), a model-agnostic framework for the probabilistic certification of LLM reasoning. Theoretically, ARES offers a novel probabilistic approach to inductively assess reasoning soundness by considering only previously validated claims, mirroring human-like error checking that discards incorrect intermediate steps. Experimentally, ARES demonstrates superior performance in robustly identifying errors in lengthy and complex reasoning chains, outperforming existing methods that degrade under error propagation.
@@
ARES’s performance is tied to the calibration of its entailment model; poor calibration can lead to unreliable scores. However, our model-agnostic approach allows for easily substituting better-calibrated components via techniques like temperature scaling to improve performance.
@@
Finally, our evaluation on four datasets with two models demonstrates effectiveness across different domains, but it is not exhaustive. Performance could also be enhanced with better prompt engineering.
@@
Acknowledgment
We thank Mayank Keoliya for discussion about this work. This research was partially supported by a gift from AWS AI to Penn Engineering’s ASSET Center for Trustworthy AI, by ASSET Center Seed Grant, ARPA-H program on Safe and Explainable AI under the award D24AC00253-00, by NSF award CCF 2442421, and by funding from the Defense Advanced Research Projects Agency’s (DARPA) SciFy program (Agreement No. HR00112520300). The views expressed are those of the author and do not reflect the official policy or position of the Department of Defense or the U.S. Government.
@@ Appendix B Method
Table A4 shows that only ARES satisfies all desiderata while none of the baseline methods does.
@@ C.1 Entailment Model
We instantiate the entailment model by prompting LLMs to judge the entailment of a hypothesis given a premise, where there can be multiple claims in the premise. The LLM’s output is either YES/NO in the binary case, or a 7-point Likert scale converted to a real value between 0 and 1.
@@ C.2 Hyperparameters for ARES
In our experiments, we used and for ARES, which determines the number of samples to take. We use for the inclusion rate for base claims to allow buffer for information overload.
@@ C.3 Experiment Details
We use a subset of examples for each experiment. Experiment results are computed using 5-fold cross-validation. For each split, the thresholds are picked for the best Macro-F1 on the validation split, and the final numbers are on the test split, averaged over the 5 folds.
C.4 Controllable Datasets
@@
One is ClaimTrees, a synthetic dataset in which the reasoning chain reasons starts from a state A, and reason all the way to another state, say T. All the reasoning rules are provided in the premise, except one, so that from that point on we know that all the claims are unsound: An example of a chain of reasoning is shown in Figure A6. In this example, rule B -> C does not actually exist, and thus the reasoning steps starting from the second derived step are unsound claims. We can construct reasoning chains with arbitrary length and errors occurring at different places.
@@
An example of results for CaptainCookRecipes is shown in Table A5. With propagated errors present, only ARES is able to capture all errors.
C.5 Computing Resources
@@
C.6 Probabilistic Entailment Model Output
To obtain probabilistic entailment model output, we instruct LLM to output one of the following: Very Likely, Likely, Somewhat Likely, Neutral, Somewhat Unlikely, Unlikely, Very Unlikely and convert them to 1, 0.8, 0.6, 0.5, 0.4, 0.2, 0.0, respectively.
@@ C.7 Best-of-N Results
For best-of-N result with standard deviations, see Table A6.
C.8 ARES Also Improves PRMs
@@
The results in Table A7 show that, while the specialized PRM is a strong baseline on its in-domain dataset (PRMBench), applying ARES significantly improves performance on the abstract ClaimTrees dataset which has many propagated errors. On out-of-domain (non-math) CaptainCook4D, ARES achieves on par performance with PRM. This demonstrates ARES’s value as a flexible, general-purpose framework that adds robustness, especially on tasks with propagated errors.
@@
| Method | Using Step Average (accstd) | Using Final Step (accstd) |
| ARES | 0.7300.045 | 0.6600.049 |
| Entail-Prev | 0.7900.043 | 0.2400.042 |
| Entail-Base | 0.5400.049 | 0.3000.046 |
| ROSCOE-LI-Self | 0.5400.051 | 0.2100.041 |
| ROSCOE-LI-Source | 0.6300.049 | 0.3100.043 |
| ReCEval-Intra | 0.4800.050 | 0.0600.024 |
| ReCEval-Inter | 0.4800.048 | 0.1900.038 |
| LLM-Judge | 0.5700.050 | 0.2500.044 |
| Dataset / Method | Qwen2.5-Math-PRM-7B | ||
| Recall | Precision | F1 | |
| PRMBench | |||
| ARES | 0.751 0.017 | 0.733 0.020 | 0.736 0.014 |
| Entail-Prev | 0.751 0.016 | 0.733 0.020 | 0.736 0.013 |
| Entail-Base | 0.643 0.022 | 0.632 0.024 | 0.624 0.018 |
| ROSCOE-LI-Self | 0.651 0.013 | 0.598 0.013 | 0.592 0.006 |
| ROSCOE-LI-Source | 0.670 0.020 | 0.621 0.019 | 0.623 0.013 |
| ReCEval-Inter | 0.644 0.014 | 0.597 0.013 | 0.596 0.009 |
| PRM | 0.763 0.020 | 0.743 0.017 | 0.749 0.016 |
| ClaimTrees-10 | |||
| ARES | 0.739 0.013 | 0.743 0.012 | 0.733 0.010 |
| Entail-Prev | 0.722 0.016 | 0.725 0.017 | 0.715 0.011 |
| Entail-Base | 0.611 0.013 | 0.616 0.013 | 0.597 0.017 |
| ROSCOE-LI-Self | 0.655 0.005 | 0.662 0.005 | 0.644 0.008 |
| ROSCOE-LI-Source | 0.604 0.020 | 0.612 0.020 | 0.591 0.024 |
| ReCEval-Inter | 0.629 0.020 | 0.628 0.019 | 0.624 0.020 |
| PRM | 0.607 0.012 | 0.622 0.013 | 0.594 0.017 |
| CaptainCook4D | |||
| ARES | 0.551 0.012 | 0.556 0.014 | 0.543 0.012 |
| Entail-Prev | 0.553 0.011 | 0.560 0.014 | 0.546 0.010 |
| Entail-Base | 0.531 0.016 | 0.533 0.017 | 0.519 0.014 |
| ROSCOE-LI-Self | 0.546 0.008 | 0.563 0.016 | 0.529 0.008 |
| ROSCOE-LI-Source | 0.469 0.015 | 0.464 0.018 | 0.457 0.017 |
| ReCEval-Inter | 0.469 0.015 | 0.465 0.018 | 0.461 0.017 |
| PRM | 0.560 0.013 | 0.569 0.017 | 0.552 0.013 |
| Dataset / Method | Recall | Precision | F1 |
| ClaimTrees-5 | |||
| ARES-1 | 0.881 | 0.900 | 0.873 |
| ARES-0.95 | 0.861 | 0.889 | 0.854 |
| ARES-bin-1 | 0.898 | 0.913 | 0.891 |
| ARES-bin-0.95 | 0.909 | 0.919 | 0.902 |
| Entail-Prev | 0.704 | 0.813 | 0.673 |
| Entail-Base | 0.830 | 0.832 | 0.824 |
| ROSCOE-LI-Self | 0.499 | 0.500 | 0.351 |
| ROSCOE-LI-Source | 0.647 | 0.650 | 0.640 |
| ReCEval-Intra | 0.500 | 0.250 | 0.332 |
| ReCEval-Inter | 0.645 | 0.648 | 0.638 |
| LLM-Judge | 0.811 | 0.864 | 0.803 |
| ClaimTrees-10 | |||
| ARES-1 | 0.937 | 0.943 | 0.936 |
| ARES-0.95 | 0.931 | 0.936 | 0.931 |
| ARES-bin-1 | 0.960 | 0.965 | 0.962 |
| ARES-bin-0.95 | 0.947 | 0.951 | 0.948 |
| Entail-Prev | 0.608 | 0.783 | 0.538 |
| Entail-Base | 0.626 | 0.636 | 0.616 |
| ROSCOE-LI-Self | 0.524 | 0.589 | 0.420 |
| ROSCOE-LI-Source | 0.544 | 0.548 | 0.533 |
| ReCEval-Intra | 0.500 | 0.247 | 0.330 |
| ReCEval-Inter | 0.566 | 0.573 | 0.555 |
| LLM-Judge | 0.767 | 0.839 | 0.750 |
| ClaimTrees-20 | |||
| ARES-1 | 0.979 | 0.979 | 0.978 |
| ARES-0.95 | 0.971 | 0.971 | 0.971 |
| ARES-bin-1 | 0.964 | 0.966 | 0.963 |
| ARES-bin-0.95 | 0.968 | 0.970 | 0.968 |
| Entail-Prev | 0.551 | 0.760 | 0.440 |
| Entail-Base | 0.533 | 0.537 | 0.522 |
| ROSCOE-LI-Self | 0.521 | 0.580 | 0.414 |
| ROSCOE-LI-Source | 0.508 | 0.509 | 0.480 |
| ReCEval-Intra | 0.500 | 0.248 | 0.331 |
| ReCEval-Inter | 0.513 | 0.516 | 0.482 |
| LLM-Judge | 0.640 | 0.788 | 0.586 |
| ClaimTrees-30 | |||
| ARES-1 | 0.973 | 0.972 | 0.971 |
| ARES-0.95 | 0.931 | 0.934 | 0.929 |
| ARES-bin-1 | 0.967 | 0.973 | 0.969 |
| ARES-bin-0.95 | 0.957 | 0.960 | 0.956 |
| Entail-Prev | 0.530 | 0.731 | 0.387 |
| Entail-Base | 0.531 | 0.539 | 0.499 |
| ROSCOE-LI-Self | 0.543 | 0.595 | 0.460 |
| ROSCOE-LI-Source | 0.498 | 0.498 | 0.461 |
| ReCEval-Intra | 0.500 | 0.262 | 0.343 |
| ReCEval-Inter | 0.506 | 0.509 | 0.464 |
| LLM-Judge | 0.581 | 0.757 | 0.482 |
| ClaimTrees-50 | |||
| ARES-1 | 0.895 | 0.899 | 0.890 |
| ARES-0.95 | 0.871 | 0.871 | 0.867 |
| ARES-bin-1 | 0.887 | 0.904 | 0.886 |
| ARES-bin-0.95 | 0.892 | 0.892 | 0.888 |
| Entail-Prev | 0.512 | 0.601 | 0.340 |
| Entail-Base | 0.507 | 0.508 | 0.486 |
| ROSCOE-LI-Self | 0.555 | 0.581 | 0.504 |
| ROSCOE-LI-Source | 0.505 | 0.509 | 0.442 |
| ReCEval-Intra | 0.500 | 0.262 | 0.343 |
| ReCEval-Inter | 0.498 | 0.496 | 0.428 |
| LLM-Judge | 0.529 | 0.714 | 0.385 |

### v2

@@ 4.4 RQ4: Is ARES useful for selecting Best-of-N generations?
Results in Table˜4 show that when using the score of the final step—a stricter and often more decisive measure—ARES significantly outperforms all other methods. Notably, the performance of simpler methods like Entail-Prev collapses on this stricter metric. This highlights ARES’s unique strength in maintaining a sound evaluation of steps throughout the entire reasoning chain, making its final assessment particularly reliable. Therefore, ARES is a strong and robust predictor of downstream task performance.
4.5 Ablations
We conduct ablations on ClaimTrees to examine the robustness and design choices of ARES.
Robustness to Irrelevant Claims.
We tested our method on reasoning trees with varying widths (irrelevant sources) and depths (path lengths), where an error was introduced by removing a single rule. As shown in Table˜A9, ARES remains stable across all configurations. In contrast, baseline methods degrade, suffering more from increased depth than width. ARES is thus capable of filtering irrelevant claims, and error propagation in long chains is the primary reason other approaches fail.
Base Claim Inclusion Probability.
We also vary the probability of including base claims and compare probabilistic vs. binary entailment models. Results in Table˜A11 show that with a probabilistic model consistently performs best, while binary models sometimes benefit from . Choosing is therefore often both accurate and efficient, as it avoids resampling base claims and reduces variance. However, as ClaimTrees has a clearer cut in soundness, the case can be different when entailment contains ambiguity.
Additional Results.
We further examine benign errors (inserted rules that do not affect downstream steps) in Section˜C.8. All methods perform equally well, unlike the irrelevant-claim setting where baselines degrade.
@@ 4.6 Discussion of Errors
Our inspection of the data and error detection outputs reveals some insights. Entail-Base fails on PRMBench because judging entailment in long math derivations is challenging. Both LLM-Judge and Entail-Base fail in DeltaBench, with Entail-Base struggling to judge entailment in very long reasoning chains. In naturally occurring datasets, error propagation is limited and not always annotated, so Entail-Prev performs close to ARES. However, synthetic data shows Entail-Prev fails with propagated errors. LLM-Judge sometimes fails to follow instructions, outputting incorrect numbers of scores relative to claims being judged. Pairwise methods in ROSCOE and ReCEval cannot detect complex errors that need multiple claims as premise. ARES can only improve upon entailment models that can already do correct entailment.
@@ 5 Related Work
Guarantees for Single-Step Explanations.
Research in interpretability has shifted from heuristic evaluation toward formal guarantees for individual predictions. One major branch is inherently interpretable models, which provide guarantees such as optimality (Angelino et al., 2018; Ustun and Rudin, 2019), monotonicity (Gupta et al., 2016; Milani Fard et al., 2016), or faithfulness by construction in deep learning models (You et al., 2025; Bassan et al., 2025b). A second branch focuses on post-hoc explanations for black-box models, including conservation guarantees (Bach et al., 2015; Shrikumar et al., 2017; Montavon et al., 2017), local accuracy, missingness, and consistency (Lundberg and Lee, 2017; Wu et al., 2024b), precision (Ribeiro et al., 2018), minimality (Ferreira et al., 2022; Bassan and Katz, 2023; Bassan et al., 2025b), sufficiency (Bassan et al., 2025b, a), or certified interventions via recourse methods (Ustun et al., 2019; Karimi et al., 2020). A third branch of work provides certified robustness guarantees, particularly in the form of stability certificates (Xue et al., 2023; Kim et al., 2024; Jin et al., 2025), which have been applied to model explainability in medicine (Achara et al., 2025). Our work extends stability guarantees to LLM reasoning chains.
Guarantees and Verification for Multi-Step Reasoning.
While single-step methods are well-studied, LLMs often generate multi-step reasoning chains prone to hallucinations and error propagation (Huang et al., 2025; Lyu et al., 2024). A significant body of work focuses on practical error detection without formal guarantees, including self-consistency checkers (Manakul et al., 2023; Dhuliawala et al., 2024) and automated verifiers such as LLM Judges (Tyagi et al., 2024; He et al., 2024, 2025), Process Reward Models (PRMs) (Lightman et al., 2024; Zhang et al., 2025), and specialized entailment models (Dalvi et al., 2021; Weir et al., 2024; Havaldar et al., 2025).
To provide more rigor, logic-based verifiers assess soundness, though often limiting to pairwise checks (Golovneva et al., 2023; Prasad et al., 2023) or taking a brittle approach to propagated errors (Mukherjee et al., 2025). An early formal guarantee, Faithful Chain-of-Thought, ensures reasoning traces deterministically yield the final answer (Lyu et al., 2023). While a suite of evaluation benchmarks exists (Tyagi et al., 2024; Jacovi et al., 2024; Song et al., 2025; Zheng et al., 2025; He et al., 2025), a unified standard for error definition is still emerging (Lee and Hockenmaier, 2025; Mukherjee et al., 2025). Recent statistical methods provide calibrated step-level reliability for generation but focus on isolated predictions within the chain (Feng et al., 2025; Quach et al., 2024; Cherian et al., 2024). In contrast, our work introduces propagation-aware guarantees that certify entire reasoning chains, ensuring upstream errors do not corrupt downstream judgments.
@@ 6 Conclusion
Current methods cannot reliably detect propagation errors in LLM reasoning chains. To address this limitation, we introduce ARES, a novel framework for certifying the soundness of an LLM’s reasoning chain. By quantifying the soundness of each claim through autoregressive sampling, ARES provides a fine-grained inductive guarantee on the chain’s overall reliability that is useful in error detection. Empirically, ARES demonstrates superior performance, robustly identifying errors in lengthy and complex reasoning chains where existing methods fail due to error propagation.
@@
ARES’s performance is tied to the quality of the entailment model; poor calibration can lead to unreliable scores. However, our model-agnostic approach allows for easily substituting better-calibrated components via techniques like temperature scaling to improve performance.
@@
Finally, our evaluation on four datasets with two models demonstrates effectiveness across different domains, but it is not exhaustive. Performance could also be improved with better LLM prompts.
@@
Acknowledgments
We thank Mayank Keoliya for discussion about this work. This research was partially supported by a gift from AWS AI to Penn Engineering’s ASSET Center for Trustworthy AI, by ASSET Center Seed Grant, ARPA-H program on Safe and Explainable AI under the award D24AC00253-00, by NSF award CCF 2442421, by the AI2050 program at Schmidt Sciences (Grant G-25-67983), and by funding from the Defense Advanced Research Projects Agency’s (DARPA) SciFy program (Agreement No. HR00112520300). The views expressed are those of the author and do not reflect the official policy or position of the Department of Defense or the U.S. Government.
@@ Appendix B Method
Table˜A5 shows that only ARES satisfies all desiderata while none of the baseline methods does.
@@ C.1 Entailment Model
We instantiate an LLM judge to assess whether a hypothesis is supported by a premise that may contain multiple claims. We use two output modes: (i) binary YES/NO, mapped to ; and (ii) a 7-point Likert scale where the LLM must output exactly one label from {Very Likely, Likely, Somewhat Likely, Neutral, Somewhat Unlikely, Unlikely, Very Unlikely}. We convert the label to a probability via
| Very Likely | |||
| Likely | |||
| Somewhat Likely | |||
| Neutral | |||
| Somewhat Unlikely | |||
| Unlikely | |||
| Very Unlikely |
which fits in a double-column layout.
Contradiction scoring. For contradiction judgments (e.g., in ROSCOE and ReCEVal), we apply the same labels but invert the mapping so that higher scores indicate stronger contradiction:
| Very Unlikely | |||
| Unlikely | |||
| Somewhat Unlikely | |||
| Neutral | |||
| Somewhat Likely | |||
| Likely | |||
| Very Likely |
with the binary case mapped analogously (YES/NO for “is contradiction?”).
@@ C.2 Hyperparameters for ARES
In our experiments, we used and for ARES, which determines the number of samples to take. We use for the inclusion rate for base claims to allow buffer for information overload. The hyperparameters and are chosen following previous work (Jin et al., 2025). The prompts are tuned by manually examining the results and seeing that they are able to produce reasonable results for entailment on each dataset. We observe that DeltaBench prefers simpler and more natural prompts while for other datasets the prompts need to have more specifications for what entail vs. not entail mean. The prompts are released in the codebase https://github.com/fallcat/ares.
@@ C.3 Experiment Details
We use a subset of examples for each experiment. Experiment results are computed using 5-fold cross-validation. For each split, the thresholds are picked for the best Macro-F1 on the validation split, and the final numbers are on the test split, averaged over the 5 folds. The standard deviation is reported for the 5 folds. Specific packages used can be found in our codebase https://github.com/fallcat/ares.
C.4 Licenses for Artifacts
All datasets, models, and code used in this work follow the licenses and terms specified by their original authors, as cited in the corresponding papers. We do not redistribute these artifacts under different terms. For the artifacts we release (https://github.com/fallcat/ares) code and evaluation scripts), we provide them under the MIT license.
Our use of existing artifacts is consistent with their intended use as specified by the original authors (e.g., datasets accessed for research purposes were used only in research contexts). For the artifacts we create, we specify that they can be freely used, modified, and redistributed under the MIT license.
C.5 Controllable Datasets
| Method | Using Step Average (accstd) | Using Final Step (accstd) |
| ARES | 0.7300.045 | 0.6600.049 |
| Entail-Prev | 0.7900.043 | 0.2400.042 |
| Entail-Base | 0.5400.049 | 0.3000.046 |
| ROSCOE-LI-Self | 0.5400.051 | 0.2100.041 |
| ROSCOE-LI-Source | 0.6300.049 | 0.3100.043 |
| ReCEval-Intra | 0.4800.050 | 0.0600.024 |
| ReCEval-Inter | 0.4800.048 | 0.1900.038 |
| LLM-Judge | 0.5700.050 | 0.2500.044 |
@@ C.5 Controllable Datasets
ClaimTrees is a synthetic dataset in which the reasoning chain reasons starts from a state such as AZ, and reason all the way to another state, say VD. All the reasoning rules are provided in the premise, except one, so that from that point on we know that all the claims are unsound: An example of a chain of reasoning is shown in Figure˜A6. In this example, rule H3 -> VD does not actually exist, and thus the reasoning steps starting from the third derived step onward are unsound claims. We can construct reasoning chains with arbitrary length and errors occurring at different places.
@@
An example of results for CaptainCookRecipes is shown in Table˜A8. With propagated errors present, only ARES is able to capture all errors.
C.6 Computing Resources
@@ C.7 Best-of-N Results
For best-of-N result with standard deviations, see Table˜A6.
C.8 Ablations
| Dataset / Method | Qwen2.5-Math-PRM-7B | ||
| Recall | Precision | F1 | |
| PRMBench | |||
| ARES | 0.751 0.017 | 0.733 0.020 | 0.736 0.014 |
| Entail-Prev | 0.751 0.016 | 0.733 0.020 | 0.736 0.013 |
| Entail-Base | 0.643 0.022 | 0.632 0.024 | 0.624 0.018 |
| ROSCOE-LI-Self | 0.651 0.013 | 0.598 0.013 | 0.592 0.006 |
| ROSCOE-LI-Source | 0.670 0.020 | 0.621 0.019 | 0.623 0.013 |
| ReCEval-Inter | 0.644 0.014 | 0.597 0.013 | 0.596 0.009 |
| PRM | 0.763 0.020 | 0.743 0.017 | 0.749 0.016 |
| ClaimTrees-10 | |||
| ARES | 0.739 0.013 | 0.743 0.012 | 0.733 0.010 |
| Entail-Prev | 0.722 0.016 | 0.725 0.017 | 0.715 0.011 |
| Entail-Base | 0.611 0.013 | 0.616 0.013 | 0.597 0.017 |
| ROSCOE-LI-Self | 0.655 0.005 | 0.662 0.005 | 0.644 0.008 |
| ROSCOE-LI-Source | 0.604 0.020 | 0.612 0.020 | 0.591 0.024 |
| ReCEval-Inter | 0.629 0.020 | 0.628 0.019 | 0.624 0.020 |
| PRM | 0.607 0.012 | 0.622 0.013 | 0.594 0.017 |
| CaptainCook4D | |||
| ARES | 0.551 0.012 | 0.556 0.014 | 0.543 0.012 |
| Entail-Prev | 0.553 0.011 | 0.560 0.014 | 0.546 0.010 |
| Entail-Base | 0.531 0.016 | 0.533 0.017 | 0.519 0.014 |
| ROSCOE-LI-Self | 0.546 0.008 | 0.563 0.016 | 0.529 0.008 |
| ROSCOE-LI-Source | 0.469 0.015 | 0.464 0.018 | 0.457 0.017 |
| ReCEval-Inter | 0.469 0.015 | 0.465 0.018 | 0.461 0.017 |
| PRM | 0.560 0.013 | 0.569 0.017 | 0.552 0.013 |
Inserting Irrelevant Claims.
Table˜A9 shows that for reasoning chains about trees of depth 5 with 3 sources, other methods show more significant performance drops while ARES maintains high performance. For wide reasoning trees, other methods don’t drop as much, but in trees with greater depth, their error rates increase significantly while ARES remains stable.
Inserting Benign Errors.
We also insert non-existing rules which don’t affect later steps into reasoning chains. Table A10 shows that all methods perform equally well when errors don’t cause downstream propagation.
We allow using different for the flexibility to not include all the base claims in the premise, and we want to see the impact of different design choices. Table˜A11 shows ablations for using vs. as well as using granular vs binary entailment models (which use strict ). consistently performs better with a probabilistic entailment model, while a binary entailment model sometimes benefits from . Thus we can effectively choose , including all base claims, which can greatly reduce the actual computation cost as the effective sample size now only depends on entailment of derived claims.
C.9 ARES Also Improves PRMs
@@
The results in Table˜A7 show that, while the specialized PRM is a strong baseline on its in-domain dataset (PRMBench), applying ARES significantly improves performance on the abstract ClaimTrees dataset which has many propagated errors. On out-of-domain (non-math) CaptainCook4D, ARES achieves on par performance with PRM. This demonstrates ARES’s value as a flexible, general-purpose framework that adds robustness, especially on tasks with propagated errors.
Appendix D AI Assistants in Research
We used AI assistants (e.g., ChatGPT, Cursor) to support coding frameworks, generate visualizations, and revise writing for clarity and readability.
@@ Appendix D AI Assistants in Research
| Dataset / Method | Recall | Precision | F1 |
| ClaimTrees-5 | |||
| ARES-1 | 0.881 | 0.900 | 0.873 |
| ARES-0.95 | 0.861 | 0.889 | 0.854 |
| ARES-bin-1 | 0.898 | 0.913 | 0.891 |
| ARES-bin-0.95 | 0.909 | 0.919 | 0.902 |
| Entail-Prev | 0.704 | 0.813 | 0.673 |
| Entail-Base | 0.830 | 0.832 | 0.824 |
| ROSCOE-LI-Self | 0.499 | 0.500 | 0.351 |
| ROSCOE-LI-Source | 0.647 | 0.650 | 0.640 |
| ReCEval-Intra | 0.500 | 0.250 | 0.332 |
| ReCEval-Inter | 0.645 | 0.648 | 0.638 |
| LLM-Judge | 0.811 | 0.864 | 0.803 |
| ClaimTrees-10 | |||
| ARES-1 | 0.937 | 0.943 | 0.936 |
| ARES-0.95 | 0.931 | 0.936 | 0.931 |
| ARES-bin-1 | 0.960 | 0.965 | 0.962 |
| ARES-bin-0.95 | 0.947 | 0.951 | 0.948 |
| Entail-Prev | 0.608 | 0.783 | 0.538 |
| Entail-Base | 0.626 | 0.636 | 0.616 |
| ROSCOE-LI-Self | 0.524 | 0.589 | 0.420 |
| ROSCOE-LI-Source | 0.544 | 0.548 | 0.533 |
| ReCEval-Intra | 0.500 | 0.247 | 0.330 |
| ReCEval-Inter | 0.566 | 0.573 | 0.555 |
| LLM-Judge | 0.767 | 0.839 | 0.750 |
| ClaimTrees-20 | |||
| ARES-1 | 0.979 | 0.979 | 0.978 |
| ARES-0.95 | 0.971 | 0.971 | 0.971 |
| ARES-bin-1 | 0.964 | 0.966 | 0.963 |
| ARES-bin-0.95 | 0.968 | 0.970 | 0.968 |
| Entail-Prev | 0.551 | 0.760 | 0.440 |
| Entail-Base | 0.533 | 0.537 | 0.522 |
| ROSCOE-LI-Self | 0.521 | 0.580 | 0.414 |
| ROSCOE-LI-Source | 0.508 | 0.509 | 0.480 |
| ReCEval-Intra | 0.500 | 0.248 | 0.331 |
| ReCEval-Inter | 0.513 | 0.516 | 0.482 |
| LLM-Judge | 0.640 | 0.788 | 0.586 |
| ClaimTrees-30 | |||
| ARES-1 | 0.973 | 0.972 | 0.971 |
| ARES-0.95 | 0.931 | 0.934 | 0.929 |
| ARES-bin-1 | 0.967 | 0.973 | 0.969 |
| ARES-bin-0.95 | 0.957 | 0.960 | 0.956 |
| Entail-Prev | 0.530 | 0.731 | 0.387 |
| Entail-Base | 0.531 | 0.539 | 0.499 |
| ROSCOE-LI-Self | 0.543 | 0.595 | 0.460 |
| ROSCOE-LI-Source | 0.498 | 0.498 | 0.461 |
| ReCEval-Intra | 0.500 | 0.262 | 0.343 |
| ReCEval-Inter | 0.506 | 0.509 | 0.464 |
| LLM-Judge | 0.581 | 0.757 | 0.482 |
| ClaimTrees-50 | |||
| ARES-1 | 0.895 | 0.899 | 0.890 |
| ARES-0.95 | 0.871 | 0.871 | 0.867 |
| ARES-bin-1 | 0.887 | 0.904 | 0.886 |
| ARES-bin-0.95 | 0.892 | 0.892 | 0.888 |
| Entail-Prev | 0.512 | 0.601 | 0.340 |
| Entail-Base | 0.507 | 0.508 | 0.486 |
| ROSCOE-LI-Self | 0.555 | 0.581 | 0.504 |
| ROSCOE-LI-Source | 0.505 | 0.509 | 0.442 |
| ReCEval-Intra | 0.500 | 0.262 | 0.343 |
| ReCEval-Inter | 0.498 | 0.496 | 0.428 |
| LLM-Judge | 0.529 | 0.714 | 0.385 |